what is the slope of the direct variation y=-2/3x?? how do you find it. what is the relationship between the constant of variation and the slope???
Answer by stanbon(75887)

what is the slope of the direct variation y=-2/3x?? how do you find it. what is the relationship between the constant of variation and the slope???
-------------
slope = [change in y]/[change in x"
When x=0 y=0
When x=1 y=-2/3
So slope = [0--2/3]/[0-1]=-2/3
Notice that is the coefficient of x.
----------------
In your equation x and y are directly related.
The constant of variation in this particular
direct relationship is -2/3 which happens to
also be the slope.
----------
Cheers,
Stan H.